# Diagnosing Vision Language Models’ Perception by Leveraging Human Methods for Color Vision Deficiencies

### v1

@@ Appendix C Language Knowledge of the Ishihara Test and CVDs
Before testing visual ability, we verified that all LVLMs possess basic language knowledge of the Ishihara Test and the three major color vision deficiencies. As the main paper includes only one example (Table 2), this appendix provides full outputs for the remaining models.
@@
Tables 1418 show representative answers from Llama-3.2, LLaVA-NeXT, mPLUG-Owl3, Phi-3.5, and Qwen2.5-VL. Each table has input prompts on the left and trimmed model responses on the right. All models—except Phi-3.5, which returned an error for the Ishihara query—accurately describe the test and define Protanopia, Deuteranopia, and Tritanopia, confirming sufficient textual knowledge. These results show that failures in the main experiments arise from weaknesses in visual grounding, not language understanding.
Appendix D Plate-level Model Outputs
The main paper reports only aggregated accuracy due to space constraints. Since the Ishihara subset includes just 25 plates, we provide the complete set of raw predictions here. Tables 11, 12, and 13 show the digits predicted by each model under the Base, Linguistic Support, and Visual Support prompts. The first row labeled Gold indicates the reference digit for each condition [^3] Plates without a digit or with “Nothing” as the correct response are marked as N/A.. Lower rows show the outputs returned by each LVLM.
Base prompt outputs
Table 11 shows model predictions and gold digits. In Protanopia and Deuteranopia, most models return digits matching the Normal condition, indicating failure to simulate red and green confusion. Failure patterns vary: Phi-3.5 produces no outputs, while Qwen2.5-VL repeatedly generates "70" or similar values, likely due to decoding issues. GPT-4o and mPLUG-Owl3 perform well on Normal and Tritanopia but diverge under red and green conditions. These results show that strong performance under normal vision does not imply accurate CVD simulation. The Base setting shows that current LVLMs often rely on shallow heuristics and fail to capture red and green perceptual variation, as reflected in Table 4.
Linguistic support outputs
Table 12 shows predictions under prompts that include brief descriptions of each condition. Despite the added context, accuracy remains similar to the Base setting. Models often respond as if assuming normal vision, suggesting that descriptive text does not influence internal visual reasoning. Qwen2.5-VL outputs under Base but fails under Linguistic Support, suggesting that disability-related terms can disrupt generation. Overall, text-only prompts neither improve grounding nor help simulate perceptual differences, showing the limitations of linguistic cues alone.
Visual support outputs
Table 13 shows predictions under Visual Support prompts, which include a reference image before the target plate. LLaVA-NeXT and Llama-3.2 are excluded due to lack of multi-image support. Evaluated models are mPLUG-Owl3, Phi-3.5, Qwen2.5-VL, and GPT-4o. No model improves over the Base setting, and many perform worse. mPLUG-Owl3 and GPT-4o often fail on plates they previously answered correctly, suggesting the added image introduces noise rather than clarity. N/A responses increase, with GPT-4o and Qwen2.5-VL more often producing empty or degenerate outputs. These results suggest that visual examples may cause confusion or overgeneralization. Overall, Visual Support fails to enhance simulation and often increases errors, showing that current LVLMs struggle to leverage visual demonstrations effectively.
Appendix E Per-token Perplexity Computation
We report here the exact procedure used to derive the confidence scores in §3.4. For each Ishihara plate, the image, prompt, and gold numeral answer are concatenated into a single input sequence. All tokens that precede the answer are masked with ignore_index, ensuring that only the answer tokens contribute to the loss. Given answer tokens $x_{1\!\,:\,\!L}$, we obtain the token-level log-probabilities:
$v_{i}=\log P(x_{i}\mid\texttt{prompt},x_{1\!\,:\,\!i-1}),$
and compute the average negative log-likelihood:
$\mathcal{L}=-\frac{1}{L}\sum_{i=1}^{L}v_{i}\quad\text{and}\quad\mathrm{PPL}=\exp(\mathcal{L}).$
Lower $\mathrm{PPL}$ thus corresponds to higher model confidence in the forced-decoded answer. All perplexity figures reported in Figure 2 and Appendix F are obtained via this definition.
Appendix F Perplexity and Model Confidence
Token-level perplexity is measured for Llama-3.2, mPLUG-Owl3, Phi-3.5, and Qwen2.5-VL. LLaVA-NeXT hides the answer during forced decoding, and GPT-4o does not return token probabilities, so both are excluded.
Base prompt
Table 19 reports mean and standard deviation of perplexity across vision types. All four models show the same trend: very low values for Normal and Tritanopia, and much higher ones for red and green deficiencies. For instance, Llama-3.2 rises from 33716.2 to 513.0, and mPLUG-Owl3 from 5.3 to 353.0. This large gap, along with low variance for Normal and high for CVDs, suggests a strong bias toward normal vision.
Linguistic Support
Table 20 shows that brief textual descriptions have little effect on the mean and often increase variance. Llama-3.2 and Phi-3.5 widen the gap between Normal and CVDs. Phi-3.5 rises from 17.4 to 171.7, while Qwen2.5-VL changes only slightly. Thus, linguistic context fails to shift the models’ perceptual prior.
Visual Support
Table 21 covers three models with multi-image input. mPLUG-Owl3 drops from 390.2 to 39.4, and Phi-3.5 from 171.7 to 11.1, but neither improves accuracy. Qwen2.5-VL becomes unstable, with Protanopia variance at 362.0. Little change in means and higher variance suggest visual support adds noise, not clarity.
Appendix G Logit Lens Probability Computation
At each transformer layer we project the hidden state $h$ back into the vocabulary space with the unembedding matrix $W_{\text{unembed}}$ and apply a softmax:
$\text{LogitLens}(h)\;=\;\text{softmax}\bigl{(}h\cdot W_{\text{unembed}}\bigr{)}.$
The resulting distribution gives the probability of every token at that layer. For a specific diagnosis-label token $t$ (e.g., Protanopia or Deuteranopia) we use
$P(t)\;=\;\frac{\exp\!\bigl{(}(h\cdot W_{\text{unembed}})_{t}\bigr{)}}{\sum_{j}\exp\!\bigl{(}(h\cdot W_{\text{unembed}})_{j}\bigr{)}}.$
Here, $(h\cdot W_{\text{unembed}})_{t}$ denotes the logit for token $t$, and the denominator normalizes over all vocabulary tokens. This $P(t)$ is the layer-wise probability assigned to the diagnosis label.
Appendix H Detailed Diagnosis of CVDs with LVLMs
@@
Figures 710 extend the main analysis by visualising token-level label probabilities across transformer layers for four additional LVLMs. Each panel groups plates by the correct vision condition and plots plate-wise trajectories (thin lines) together with the average curve (bold line). Left plots correspond to Normal or Tritanopia, right plots to Protanopia or Deuteranopia.
@@
Llama-3.2 (Figure 7), For Normal/Tritanopia the Normal label stays near zero until three layers before the head, then rises above 50%. red and green show an analogous late surge in the Protanopia curve, mirroring the perplexity jump in Table 19.
@@
mPLUG-Owl3 (Figure 8), Early layers are mixed but converge rapidly. Tritanopia exceed 80% by layer 26. When the digit is unreadable, the model shifts probability toward Tritanopia, possibly explaining its low variance under Base.
@@
Phi-3.5 (Figure 9), Curves remain diffuse and never exceed 25% even near the output. This matches the wide perplexity spread and frequent no-answer cases in Table 11.
@@
Qwen2.5-VL (Figure 10), Normal show local peaks rather than a smooth ascent. red and green plates present competing red and green waves that collapse to a single token at the final layer. However, the overall probabilities remain extremely low, on the order of $10^{-6}$ even beyond layer 6, which indicates very weak model confidence.
@@
Table 8: Base prompts for each color-vision condition.
@@ A Tritanopic person sees this number: {few_shot_number}
Table 9: Prompts with visual support examples for each color-vision condition.
Setting Prompt
Protanopia You are Protanopic.
What number do you see?
Deuteranopia You are Deuteranopic.
What number do you see?
Tritanopia You are Tritanopic.
What number do you see?
Normal Vision What number do you see?
Table 10: Prompts with linguistic support for each color-vision condition.
Normal Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
Llama 19 8 8 42 7 1 2 4 15 2 42 42 1 7 7 4 53 15 15 3 4 52 52 52 8
Llava 12 8 6 29 37 3 6 16 7 2 6 9 5 6 7 16 3 1 1 1 1 28 42 35 96
mPlug 12 8 6 29 37 5 3 15 24 2 6 9 10 5 7 16 13 100 100 10 100 23 42 35 96
Phi
Qwen 12 8 6 29 57 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 48 74 95
GPT 12 8 6 29 57 5 3 15 74 2 6 74 45 5 7 16 73 74 74 74 5 93 42 35 96
Llama 19 2 5 17 3 3 2 6 3 2 2 2 45 7 4 6 46 73 4 53 52 4 4 4 2
Llava 12 8 6 29 57 3 6 16 7 2 6 9 1 6 7 16 3 1 1 1 1 20 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 1 1 1 1 23 42 35 96
Phi
Qwen 21 8 9 29 74 74 74 74 74 74 74 74 74 74 74 74 7 74 74 74 74 38 43 74 95
GPT 12 3 3 29 55 3 5 15 4 9 3 37 45 N/A 7 16 13 74 2 74 27 25 N/A 39 96
Llama 19 2 5 52 1 53 52 49 45 53 52 52 36 55 7 56 4 4 2 49 53 42 52 37 8
Llava 12 8 6 29 57 3 3 16 14 2 6 9 10 6 7 16 3 10 1 1 12 20 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 2 6 9 10 5 7 16 13 1 1 10 1 23 42 35 96
Phi
Qwen 21 8 9 29 74 74 74 74 74 74 74 74 74 74 74 53 7 74 74 74 74 88 43 30 95
GPT 12 8 6 29 57 3 8 N/A 74 2 6 97 45 3 7 16 73 74 N/A N/A 5 93 42 35 96
Llama 19 2 5 17 1 3 2 4 1 3 4 42 46 4 7 3 8 4 4 4 1 3 4 3 8
Llava 12 8 6 29 37 3 3 16 7 2 6 9 1 6 7 16 3 1 1 1 12 28 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 10 10 10 10 23 42 35 96
Phi
Qwen 21 8 6 29 37 74 74 74 74 74 74 74 74 5 3 53 3 74 0 74 74 38 43 30 95
GPT 12 8 6 29 57 N/A 3 15 74 2 3 37 36 5 7 16 33 74 74 6 3 N/A 42 35 96
Table 11: Predicted outputs of six LVLMs (Llama, Llava, mPlug, Phi, Qwen, and GPT) on Ishihara color-vision test plates under four simulated conditions (Normal, Protanopia, Deuteranopia, Tritanopia) using the Base prompt only. The “Gold” row shows the ground-truth digits.
@@
Table 12: Predicted outputs of six LVLMs (Llama, Llava, mPlug, Phi, Qwen, and GPT) on Ishihara color-vision test plates under four simulated conditions, using Linguistic Support prompts that include a brief explanation of how each deficiency alters color perception. The “Gold” row shows the ground-truth digits.
@@
Table 13: Predicted outputs of four LVLMs (mPlug, Phi, Qwen, and GPT) on Ishihara color-vision test plates under four simulated conditions, using Visual Support prompts that present an illustrative example image of each plate as seen under the deficiency. The “Gold” row shows the ground-truth digits. Evaluation was performed only on models capable of ingesting two or more images in a single turn.
@@
Normal Protanopia Deuteranopia Tritanopia
LVLM Size Mean SD Mean SD Mean SD Mean SD
LLaMA 11B 33716.2 176261.9 513.0 1453.1 524.7 1451.5 33716.2 176261.9
mPLUG-Owl3 7B 5.3 7.7 353.0 646.4 331.9 652.0 5.3 7.7
Phi-3.5 4.2B 17.2 25.3 119.5 406.4 88.6 404.2 17.2 25.3
Qwen2.5-VL 7B 31.4 62.2 64.5 91.2 44.9 74.4 31.4 62.2
Table 19: Mean and standard deviation (SD) of perplexity for each vision condition under the Base prompt setting.
Normal Protanopia Deuteranopia Tritanopia
LVLM Size Mean SD Mean SD Mean SD Mean SD
LLaMA 11B 14140.3 66220.6 596.8 1525.2 605.8 1523.3 14140.3 66220.6
mPLUG-Owl3 7B 5.3 7.4 309.2 621.3 288.0 626.2 5.3 7.4
Phi-3.5 4.2B 17.4 22.9 171.7 531.1 116.4 522.7 17.4 22.9
Qwen2.5-VL 7B 32.0 59.6 58.0 85.4 41.3 69.8 32.0 59.6
Table 20: Mean and standard deviation (SD) of perplexity for each vision condition under the Linguistic prompt setting.
Normal Protanopia Deuteranopia Tritanopia
LVLM Size Mean SD Mean SD Mean SD Mean SD
LLaMA 11B
mPLUG-Owl3 7B 11.0 27.3 39.4 55.7 30.8 47.3 11.0 27.3
Phi-3.5 4.2B 8.9 11.3 11.1 13.2 9.9 13.3 8.9 11.3
Qwen2.5-VL 7B 100.7 132.9 248.3 362.0 149.8 203.0 100.7 132.9
Table 21: Mean and standard deviation (SD) of perplexity for each vision condition under the Visual prompt setting.

### v2

@@ Appendix B Language Knowledge of the Ishihara Test and CVDs
Before testing visual ability, we verified that all LVLMs possess basic language knowledge of the Ishihara Test and the three major color vision deficiencies. As the main paper includes only one example, this appendix provides the full outputs for the remaining models.
@@
Tables 1923 show representative answers from Llama-3.2, LLaVA-NeXT, mPLUG-Owl3, Phi-3.5, and Qwen2.5-VL. Each table has input prompts on the left and trimmed model responses on the right. All models except Phi-3.5, which returned an error for the Ishihara query, accurately describe the test and define Protanopia, Deuteranopia, and Tritanopia, confirming sufficient textual knowledge. These results show that failures in the main experiments arise from weaknesses in visual grounding, not language understanding.
Prompts in Detail
For each color vision condition, we designed three types of prompts: Base, Linguistic Support, and Visual Support. The full prompt templates are provided in Tables 1416.
Table 14 presents the Base prompts, which specify only the condition (e.g., “You are Protanopic”) and then ask the model to report the number shown in the Ishihara plate. This represents the simplest setting, testing whether the model can directly simulate the specified condition.
[…]
Table 16 shows the Visual Support prompts. In this case, the condition is accompanied by a few shot reference example (e.g., “A Protanopic person sees this number: {few_shot_number}”) before asking the model to identify the number in a new plate. This setup tests whether visual exemplars serve as useful guidance for simulating CVD perception.
Together, these tables make explicit how the three prompt types were implemented, complementing the overview provided in Section 3.3.
Appendix C Effects of Prompt Variants on Simulating CVD Perception
Prompt phrasing can influence how LVLMs interpret images, particularly when the task requires aligning visual input with an altered perceptual perspective. To assess whether failures in our main experiments were attributable to prompt design rather than perceptual grounding, we evaluated four alternative prompt formulations for the Protanopia condition:
  • Prompt 1 (Description-Based): “What number would a Protanopic viewer likely see in this image? Only provide the predicted number.”
  • Prompt 2 (Doctor Perspective): “You are an eye doctor evaluating how different color vision types perceive this test plate. What number would a Protanopic patient report? Only output the number.”
  • Prompt 3 (Role-Play Perspective): “Answer as if you are a Protanopic viewer. Respond with the number you visually perceive. Only output the number.”
  • Prompt 4 (Step-by-Step Reasoning): “Step 1: Describe how the red and green dots differ under Protanopia. Step 2: Based on Step 1, what number appears to a Protanopic viewer? Only output Step 2.”
Table 12 reports digit accuracy under Normal and three CVD conditions for Prompts 1–3. Prompt 4 was excluded due to highly unstable behavior: models frequently produced free-form explanations, failed to emit a digit token in Step 2, or ignored the requested output format entirely.
Across all prompt variants, we observe moderate variability in the Normal condition, confirming that prompt phrasing can influence how LVLMs interpret visual input when no perceptual transformation is required. However, under Protanopia, Deuteranopia, and Tritanopia, accuracy remains low across prompts and models. Neither altering task framing (doctor vs. viewer) nor introducing explicit role-playing improved condition-appropriate responses. These results indicate that prompt formulation alone is insufficient to induce LVLMs to simulate condition-specific perceptual distortions, consistent with the main conclusion that limitations arise from insufficient visual grounding rather than linguistic control alone.
Normal Protanopia Deuteranopia Tritanopia
LVLM Size P1 P2 P3 P1 P2 P3 P1 P2 P3 P1 P2 P3
Llama-3.2 11B 9.5 4.8 9.5 17.6 11.8 0.0 11.8 11.8 11.8 4.8 4.8 4.8
LLaVA-NeXT 13B 61.9 52.4 57.1 5.9 5.9 5.9 5.9 5.9 5.9 47.6 61.9 57.1
mPLUG-Owl3 7B 71.4 76.2 66.7 5.9 5.9 5.9 5.9 5.9 5.9 66.7 71.4 66.7
Phi-3.5 4.2B 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Qwen2.5-VL 7B 28.6 28.6 38.1 5.9 0.0 0.0 0.0 0.0 0.0 14.3 19.1 28.6
Table 12: Digit accuracy (%) under three prompt variants (P1–P3) across simulated vision types. Prompt 4 is omitted due to unstable output formatting and inconsistent digit emission.
Appendix D Digit Results: Detailed Analysis
Tables 17 and 18 report the raw digit outputs that each LVLM produced for Ishihara plates under four simulated conditions (Normal, Protanopia, Deuteranopia, Tritanopia). Unless stated otherwise, plates marked as “N/A” (no correct digit) are excluded from accuracy and are discussed only for error behavior.
D.1 Linguistic Support prompts (Table 17).
Adding a short textual description yields only modest and inconsistent gains. Under Normal vision, GPT remains strong and LLaVA/mPLUG perform similarly to the Base setting, while Llama continues to miss many plates. Under simulated CVDs, condition-specific digits are seldom produced; the same plates that were difficult under Base remain challenging, and response biases (e.g., repeated 74/42) persist. Hallucinated digits also appear on N/A plates, showing that brief definitions alone are insufficient to induce condition-aware perception.
D.2 Visual Support prompts (Table 18).
Providing a few shot reference image before the target plate produces mixed results. GPT maintains strong accuracy under Normal vision but shows limited and inconsistent improvements for simulated CVDs. mPLUG performs slightly better on some plates compared to Base/Linguistic settings but still fails to reliably follow the instructed condition. Phi occasionally outputs correct answers without cross-plate consistency, while Qwen remains unstable and frequently repeats the same digits. Some regressions and hallucinated digits are also observed, indicating that single visual exemplars do not reliably induce condition-appropriate perception.
D.3 Overall Findings
Across all prompt types, LVLMs display recurring error patterns: frequent mode collapse to a small set of digits (e.g., 74, 42, 3), insensitivity to the instructed CVD condition, and hallucinated outputs on plates without valid digits. These behaviors show that neither short textual descriptions nor single visual exemplars are sufficient to ground condition-specific perception. While the models can describe CVDs accurately in text, they fail to translate these instructions into consistent digit recognition on Ishihara plates. This confirms that the main limitation lies in visual grounding rather than language knowledge or prompt design.
Appendix E Additional Perplexity Analyses
In the main text, we focused on mPLUG-Owl3 (Figure 2) because it supports multi-image input and shows the most stable perplexity across prompt settings. For completeness, we report violin plots of perplexity distributions for three additional models: LLaMA, Phi, and Qwen.
Figures 46 show perplexity measured by force decoding the Gold Answer for Ishihara digit predictions under four simulated vision conditions (Normal, Protanopia, Deuteranopia, Tritanopia) and three prompt types (Base, Linguistic, Visual).
Across these models, we observe similar tendencies as discussed in the main text: Normal vision yields consistently lower perplexity, while the CVD conditions lead to higher and more variable perplexity. Adding textual descriptions (Linguistic Support) or reference images (Visual Support) does not substantially alter these distributions. This supports the conclusion that prompt modifications are insufficient to reduce model uncertainty when simulating altered color vision.
Appendix F Detailed Diagnosis of CVDs with LVLMs
@@
Figures 710 extend the main analysis by visualizing token-level label probabilities across transformer layers for four additional LVLMs. Each panel groups plates by the correct vision condition and plots plate-wise trajectories (thin lines) together with the average curve (bold line). Left plots correspond to Normal or Tritanopia, right plots to Protanopia or Deuteranopia.
@@
Llama-3.2 (Figure 7), For Normal/Tritanopia, the Normal label stays near zero until three layers before the head, then rises above 50%. Red and green show an analogous late surge in the Protanopia curve, reflecting a rise in uncertainty.
@@
mPLUG-Owl3 (Figure 8), Early layers are mixed but converge rapidly. For instance, tritanopia exceeds 80% by layer 26. When the digit is unreadable, the model shifts probability toward Tritanopia, possibly explaining its low variance under Base.
@@
Phi-3.5 (Figure 9), Curves remain diffuse and never exceed 25% even near the output. This matches the wide perplexity spread and frequent no-answer cases.
@@
Qwen2.5-VL (Figure 10), Normal shows local peaks rather than a smooth ascent. Red and green plates present competing red and green waves that collapse to a single token at the final layer. However, the overall probabilities remain extremely low, on the order of $10^{-6}$ even beyond layer 6, indicating very weak model confidence.
Appendix G Fine-tuning and Inference Settings
Hyperparameter Value
Model Qwen/Qwen2-VL-2B-Instruct
Torch dtype bfloat16
Seed 42
Max input length 2048
Epochs 1
Learning rate 2e-4
Batch size 1
Gradient accumulation 32
Warmup ratio 0 (not used)
Gradient checkpointing True
LoRA rank ($r$) 8
LoRA alpha 16
LoRA dropout 0.05
Save steps 200
Eval strategy None
Table 13: Hyperparameters and settings used for fine-tuning. We used Transformers wolf-etal-2020-transformers, PEFT xu2023parameterefficientfinetuningmethodspretrained, and trl libraries with default configurations unless otherwise noted.
@@ Appendix G Fine-tuning and Inference Settings
Table 14: Base prompts for each color‐vision condition.
Setting Prompt
Protanopia You are Protanopic.
What number do you see?
Deuteranopia You are Deuteranopic.
What number do you see?
Tritanopia You are Tritanopic.
What number do you see?
Normal Vision What number do you see?
Table 15: Prompts with linguistic support for each color‐vision condition.
@@ A Tritanopic person sees this number: {few_shot_number}
Table 16: Prompts with visual support examples for each color‐vision condition.
@@
[…]
@@
[…]